OpenStudy (anonymous):

i need help solving this problem. A projectile of 1.0kg mass approaches a stationary body of 5.0kg at 10.0m/s and, after colliding, rebounds in the reverse direction along the same line with a speed of 5.0m/s. what is the speed of the 5.0kg body after the collision?

OpenStudy (stormfire1):

Well, this is clearly a momentum problem. First note, the collision is elastic...that is, the objects are bouncing off of eachother. Second, note that this is a one-dimensional problem and that momentum will be conserved before and after the collision. First, figure out the initial momentums of each object: P = mass * velocity object 1: 1kg * 10 m/s = 10 kg*m/s. object 2: 5kg * 0m/s = 0 kg*m/s Therefore, the total initial momentum is 10 kg*m/s + 0 kg*m/s = 10 kg*m/s Now see what you have after the collision object 1 = 1kg * -5 m/s = -5 kg*m/s object 2 = 5kg * V Since Pf must = Pi: 10kg*m/s = -5 kg*m/s + 5kg*V V is therefore 3 m/s for the 5kg object